OpenStudy (<3pandas18):

What is the equation of the circle with center (2,-5) that passes through the point (-2,10) Can anyone help?

OpenStudy (welshfella):

(x - 2)^2 + ( x + 5)^2 = r^2 - as the centre is at (2,-5) and r = radius

OpenStudy (welshfella):

now as the circle passes though the point (-2,10) we can plug in x = -2 and y = 10 into the above equation so we can now find the values of r^2:- (-2 -2)^2 + (-10 + 5)^2 = r^2 (-4)^2 + (-5)^2 = r^2 - calculate r^2 and you have your eqaution

OpenStudy (<3pandas18):

OMG thank you!

OpenStudy (welshfella):

yw

OpenStudy (<3pandas18):

@welshfella thank for helping but im still confused the equation you explained is not one of my answer choices.

OpenStudy (<3pandas18):